# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ II Definition of “minimal” symmetry enriched topological order
this forms the set of minimal SETs according to the definition of Jian et al. (2024). Now the GSD of a TO is equal to the number of distinct anyon types Kitaev (2006), which is $|\mathcal{C}|$ for bosonic TOs and $|\mathcal{C}|/2$ for fermionic TOs, where this factor of two compensates for our including $c\in\mathcal{C}$  [^5] The GSD is independent of the periodic/anti-periodic boundary condition.. Therefore achieving the smallest GSD will mean achieving the smallest number of anyons. We thus define
$mG_{\rm GSD}=\{\mathcal{C}\in{\rm SET}_{G}\mid|\mathcal{C}|\leq|\mathcal{C}^{\prime}|\ \forall\mathcal{C}^{\prime}\in{\rm SET}_{G}\},$ (3)
@@
as the set of SETs with minimal GSD on a torus. We can think of these as classes of minimal SETs. For a given $G$ there is a priori no reason for these two classes to have any relationship to one another. However, a useful heuristic is that the minimal class $mG_{\mathcal{D}}$ is more likely to contain Abelian TOs than the class $mG_{\rm GSD}$. This is illustrated in Fig. 2.
@@
We make this more precise by stating a result proven in App. A. It follows from the inequality $|\mathcal{C}|\leq\mathcal{D}_{\mathcal{C}}^{2}$, which we noted earlier. If there is a single Abelian TO in $mG_{\rm GSD}$, then all TOs in $mG_{\mathcal{D}}$ are Abelian and moreover $mG_{\mathcal{D}}\subseteq mG_{\rm GSD}$. On the other hand, there is no reason for there to be any Abelian TOs in $mG_{\rm GSD}$ even if one of the TOs in $mG_{\mathcal{D}}$ is Abelian. We give an example of this in App. A. It is thus more powerful to establish results about the minimal orders of $mG_{\rm GSD}$ being Abelian, than about those of $mG_{\mathcal{D}}$. Likewise, it is more powerful to establish results about the minimal orders of $mG_{\mathcal{D}}$ being non-Abelian.
@@
For our choice of $G$, specified in the next section, we will see that all TOs in $mG_{\rm GSD}$ are Abelian and thus that $mG_{\rm GSD}=mG_{\mathcal{D}}$. Thus in this case the two classifications of minimal order turn out to be equivalent. However, it is nonetheless more powerful to work with $mG_{\rm GSD}$ since this will imply things about $mG_{\mathcal{D}}$, but not the other way around. We thus henceforth use the term minimal to refer to a TO in $mG_{\rm GSD}$, unless otherwise specified.
@@ III Minimal order at fractional fillings
We consider a general system of spinless fermions that respect $\mathrm{U}(1)$ fermion number conservation and are at a fractional filling $\nu=1/q$ per unit cell; all of our arguments also easily generalize to the case $\nu=p/q$ with $p$ relatively prime to $q$. For our purposes we will henceforth take the symmetry group to be given by:
@@
where the first $\mathbb{Z}^{2}$ is translation in the $x$ and $y$ directions, the $\mathrm{U}(1)_{f}$ is $\mathrm{U}(1)$ charge with the restriction that the charge mod $2$ is equal to the fermion parity, and $\mathbb{Z}_{2}^{T}$ is the order two time-reversal operation which does not change the $\mathrm{U}(1)$ charge. In what follows we will also take the electron to be a Kramers singlet, i.e. we will assume that $\mathcal{T}^{2}=1$. This is appropriate in thinking about the possibility of the fractionalized states in TMD moiré heterostructures, where, as we emphasized, the electron spin exchange is very small compared to other energy scales. Thus the spin may be easily fully polarized in a small magnetic field without affecting the orbital motion.
@@
With this symmetry group, we demonstrate in Sec. III.2 that the minimal TOs in both $mH_{\mathcal{D}}$ and $mH_{\rm GSD}$ are Abelian with $2q$ anyons when $q$ is odd, and $4q$ anyons when $q$ is even. The proof is relatively straightforward and illustrative of how we will tackle the full $G$.
Having warmed up we dive into a discussion on constraints on the minimal TOs consistent with the full $G$. In Sec. III.3 we construct a fermionic Abelian TO which preserves $G$ and has a GSD of $4q^{2}$ ($q^{2}$) for $q$ even (odd). We then prove a number of results. First, in Sec. III.4, we show that any TO at filling $\nu=1/q$ which respects $G$ and has a GSD equal to or smaller than our construction must be Abelian. Moreover, we show that all such possibilities must have exactly the same GSD as our construction, so our construction is minimal. Then, in Sec. III.5, we show that our construction is the unique minimal TO that obey all the requirements.
@@
  • If our GSD is even smaller than this, then we can conclude that translation symmetry must be spontaneously broken. This will allow us to have a GSD as small as $q$, e.g. in a stripe order. Note that the GSD of $q$ in this case is not topological but arises from the breaking of translation symmetry.
@@ III.1 Generalities about TOs at fractional filling
In any gapped TO, bosonic or fermionic, with $\mathrm{U}(1)$ symmetry, there exists an anyon $v$ called a vison [^6] The name ‘vison’ is used here as a generalization of its original use Senthil and Fisher (2000) to denote an Ising vortex in a $\mathbb{Z}_{2}$ gauge theory. In the special situation in which the low energy physics is described by a $\mathbb{Z}_{2}$ gauge theory with a charge-$1/2$ (under the global $\mathrm{U}(1)$) chargon, and a charge-neutral vison, the latter is indeed nucleated by such a flux insertion. which is nucleated by flux insertion, whose braiding with other anyons detects their (fractional) $\mathrm{U}(1)$ charge:
$\theta_{v,a}=2\pi Q_{a},$ (6)
where $Q_{a}$ is the charge carried by $a$ mod 1. Physically, $v$ can be understood as resulting from the adiabatic insertion of a $2\pi$ $\mathrm{U}(1)$ flux. It is important to keep in mind that in a fermionic system with $\mathrm{U}(1)_{f}$ symmetry, all bosonic/fermionic local excitations carry even/odd charges.
In addition, the classic Laughlin argument shows that $Q_{v}=\sigma_{H}$ (the Hall conductance in units where electronic charge and Planck’s constant are set to 1). It follows from Eqn.  6 that $\theta_{v,v}=Q_{v}=2\pi\sigma_{H}$, and therefore that $v$ has topological spin $\theta_{v}=\pi\sigma_{H}$.
@@
We note that the background anyon $a$ must be Abelian. This was shown more formally in Cheng et al. (2016), but we sketch two complementary arguments here. First suppose that $a$ is non-Abelian, with quantum dimension $d_{a}>1$. Then the Hilbert space dimension associated with having non-Abelian anyons on every site will scale as $\sim d_{a}^{N}$, where $N$ is the number of sites Kitaev (2006). This macroscopic degeneracy must be lifted by some interaction, meaning that this description is at best useful at an intermediate energy scale, and does not correspond to the true infrared (IR) limit. Flowing towards the true IR, any further macroscopic degeneracy will continue to be broken until we have fully broken it, at which point we must have a background anyon with $d_{a}=1$, i.e. $a$ must be Abelian.
Another complementary perspective can show the same conclusion. Operationally, the background anyon is defined by the following thought experiment: 1) excite some anyon pair $x$ and $\overline{x}$, 2) adiabatically transport $x$ around a unit cell, 3) and annihilate the pair. This will return the system to its ground state, so it can only lead to a global phase $\varphi_{x}$ for the wavefunction. We can think of this phase as the braiding of $x$ around the background anyon $a$. Importantly, this phase is additive under fusion: if $z\in x\times y$, then $\varphi_{z}=\varphi_{x}+\varphi_{y}$. Via a result in Barkeshli et al. (2019) (see Eq. 45), this is then enough to establish that $\varphi_{x}=\theta_{a,x}$ for some Abelian anyon $a$, which we identify as the background anyon.
@@ III.2 Time-reversal broken minimal order
For bosons at $1/q$ filling, with $q$ even, $\mathrm{U}(1)_{q}$ (the TO of the bosonic $1/q$ Laughlin state) is a natural candidate for time-reversal breaking TOs in bosonic QCLs. We now demonstrate that for $q$ even, $\mathrm{U}(1)_{q}\boxtimes\{1,c\}$ cannot be a fermionic TO at filling $1/q$. First we note that for even $q$ $\mathrm{U}(1)_{q}$ by itself is a bosonic TO, and local operators in the theory are all bosons. It follows that the local excitation with minimal non-zero charge in $\mathrm{U}(1)_{q}$ is a charge-$2$ Cooper pair. Since every anyon in this theory has order $q$, the minimal fractional charge in this theory is $2/q$. But this means this theory does not have a background anyon of charge $Q_{a}=\nu=1/q$, contradicting the presence of $\mathrm{U}(1)_{f}$ symmetry and the filling constraint.
We can construct a possible fermionic TO by pairing two fermions into Cooper pairs. The Cooper pairs are bosons at a filling $\nu/2=1/2q$ per unit cell. We then make a TO, $\mathcal{C}$, which is translation invariant by considering $\mathrm{U}(1)_{2q}$, the bosonic Laughlin state of the Cooper pairs. Strictly speaking we have $\mathrm{U}(1)_{2q}\boxtimes\{1,c\}$, where $c$ is the fundamental fermion. This will be a fermionic TO in SETH, and is depicted in Fig. LABEL:fig:cooper. This theory has a GSD of $2q$. Of course if $q$ is odd the theory $\mathrm{U}(1)_{q}$ is a fermionic TO in SETH with a smaller GSD, $|\mathrm{U}(1)_{q}|/2=q$.
We now prove that these two constructions belong to $mH_{\rm GSD}$ for the respective choice of $q$. We will do this by showing that the Abelian sector in any fermionic TO in SETH is at least as large as our constructions. So then if our theory is to be minimal according to $mH_{\rm GSD}$ it must be Abelian with exactly the same anyon count as our constructions. This will prove that every TO in $mH_{\rm GSD}$ is Abelian and thus that $mH_{\rm GSD}=mH_{\mathcal{D}}$, so the two definitions of minimality coincide. Our proof strategy is illustrated in Fig. 4.
@@
Though we have determined the possible minimum TOs, we can still ask whether there are distinct symmetry enrichments by $\mathbb{Z}^{2}\times\mathrm{U}(1)_{f}$. To that end, we need to specify the vison anyon. It is easy to see that $v=a^{k}$, where $k$ is such that $2nk=1\pmod{q}$. Since $n$ is coprime with $q$ and $q$ is odd, such $k$ always exists and is uniquely fixed. Thus the possible minimal orders in the time reversal breaking case have Hall conductances of $\sigma_{H}=k/q\pmod{1}$ where $k$ can take all values relatively prime to $q$. It is worth noting that we can (trivially) modify the TO by stacking copies of invertible topological states [^7] In two dimensions, the basic such state is the so-called $E_{8}$ state with chiral central charge $c=8$. of charge-neutral local bosonic excitations, which does not change the anyon content.
@@
Next we tackle $q$ even. In this case we again must have that $a^{q}\neq 1$. So then $a$ must have order a multiple of $2q$. Suppose now that $a^{q}=c$, then $a^{q}$ must braid trivially with $a$. This means that $\theta_{a,a^{q}}=2q\theta_{a}=0$. Since $q$ is even we multiply by the integer $q/2$ to see that $0=q^{2}\theta_{a}=\theta_{a^{q}}$ and thus $a^{q}$ is a boson, a contradiction with $a^{q}=c$. So $c\notin\{1,a,\ldots,a^{2q-1}\}$ and we must have that there are at least $4q$ anyons in our theory. For our theory to be minimal it must thus be Abelian with $\mathcal{A}=\langle a\rangle\boxtimes\{1,c\}$ where $a$ has order exactly $2q$. All such theories are again classified Moore and Seiberg (1989); Bonderson (2007), with $\theta_{a}=\pi n/2q$ for any $n$ such that $n$ is coprime with $q$. In particular this means $n$ must be odd. We can then choose $v=a^{k}$ where $k$ is such that $nk=2\pmod{2q}$. This means $k$ must be even, and more specifically that $k=2k^{\prime}$ with $k^{\prime}$ relatively prime to $2q$. Note $n=1$ corresponds to the $\mathrm{U}(1)_{2q}$ theory, where we can choose $v=a^{2}$. Thus when $q$ is even the minimal orders have Hall conductances $\sigma_{H}=2k^{\prime}/q\pmod{1}$, where $k^{\prime}$ can take all values relatively prime to $q$.
@@
This general proof strategy will prove useful in dealing with the full $G$. We first construct a theory in SETG, then show that all minimal theories must be at least as big as our construction. We again refer to reader to Fig. 4 for an illustration of this strategy. Having obtained the anyon count of each minimal theory we then show that in the presence of time reversal symmetry the minimal TO is in fact unique.
@@
Figure 4: Our proof strategy for demonstrating the number of anyons a minimal theory must have. Using Cooper pairing we construct a theory in SETG. This theory’s anyon count must upper bound the anyon count of any minimal SET. We further consider all possible powers of the Abelian anyons we know our theory must possess. This will lower bound the anyon count. Any putative minimal SET will be caught in the vice between these two bounds.

### v2

@@ II Definition of “minimal” symmetry enriched topological order
this forms the set of minimal SETs according to the definition of Ref. Jian et al. (2024). Now the GSD of a TO is equal to the number of distinct anyon types Kitaev (2006), which is $|\mathcal{C}|$ for bosonic TOs and $|\mathcal{C}|/2$ for fermionic TOs, where this factor of two compensates for our including $c\in\mathcal{C}$ [^9] The GSD is independent of the periodic or antiperiodic boundary condition. As mentioned before, we assume an even number of electrons. Therefore, achieving the smallest GSD will mean achieving the smallest number of anyons. We thus define
$mG^{(\nu)}_{\rm GSD}=\{\mathcal{C}\in{\rm SET}_{G,\nu}\mid|\mathcal{C}|\leq|\mathcal{C}^{\prime}|\ \forall\mathcal{C}^{\prime}\in{\rm SET}_{G,\nu}\},$ (3)
@@
as the set of SETs with minimal GSD on a torus. We can think of these as classes of minimal SETs. For a given $G$ there is a priori no reason for these two classes to have any relationship to one another. However, a useful heuristic is that the minimal class $mG^{(\nu)}_{\mathcal{D}}$ is more likely to contain Abelian TOs than the class $mG^{(\nu)}_{\rm GSD}$. This is illustrated in Fig. 2.
@@
We make this more precise by stating a result proven in Appendix A. It follows from the inequality $|\mathcal{C}|\leq\mathcal{D}_{\mathcal{C}}^{2}$, which we noted earlier. If there is an Abelian TO in $mG^{(\nu)}_{\rm GSD}$, then all TOs in $mG^{(\nu)}_{\mathcal{D}}$ are Abelian and moreover $mG^{(\nu)}_{\mathcal{D}}\subseteq mG^{(\nu)}_{\rm GSD}$. On the other hand, there is no reason for there to be any Abelian TOs in $mG^{(\nu)}_{\rm GSD}$ even if one of the TOs in $mG^{(\nu)}_{\mathcal{D}}$ is Abelian. We give an example of this in Appendix A. It is thus more powerful to establish results about the minimal orders of $mG^{(\nu)}_{\rm GSD}$ being Abelian, than about those of $mG^{(\nu)}_{\mathcal{D}}$. Likewise, it is more powerful to establish results about the minimal orders of $mG^{(\nu)}_{\mathcal{D}}$ being non-Abelian.
@@
For our choice of $G$, specified in the next section, we see that all TOs in $mG^{(\nu)}_{\rm GSD}$ are Abelian and thus that $mG^{(\nu)}_{\rm GSD}=mG^{(\nu)}_{\mathcal{D}}$. Thus in this case the two classifications of minimal order turn out to be equivalent. However, it is nonetheless more powerful to work with $mG^{(\nu)}_{\rm GSD}$ since this will imply things about $mG^{(\nu)}_{\mathcal{D}}$, but not the other way around. We thus henceforth use the term minimal to refer to a SET in $mG^{(\nu)}_{\rm GSD}$, unless otherwise specified. Note that we often discuss minimal TOs rather than minimal SETs, here a minimal TO is the underlying TO of a minimal SET.
@@ III Minimal order at fractional fillings
We consider a general system of spinless fermions that respect $\mathrm{U}(1)$ fermion number conservation and are at a fractional filling $\nu=1/q$ per unit cell; all of our arguments also easily generalize to the case $\nu=p/q$ with $p$ relatively prime to $q$. For our purposes we henceforth take the symmetry group to be given by
@@
where the first $\mathbb{Z}^{2}$ is translation in the $x$ and $y$ directions, the $\mathrm{U}(1)_{f}$ is $\mathrm{U}(1)$ charge with the restriction that the charge mod $2$ is equal to the fermion parity, and $\mathbb{Z}_{2}^{T}$ is the order-two time-reversal operation which does not change the $\mathrm{U}(1)$ charge. In what follows we will also take the electron to be a Kramers singlet i.e., we assume that $\mathcal{T}^{2}=1$. This is appropriate in thinking about the possibility of the fractionalized states in TMD moiré heterostructures, where, as we emphasized, the electron spin exchange is very small compared with other energy scales. Thus the spin may be easily fully polarized in a small magnetic field without affecting the orbital motion.
@@
We review some generalities about TOs respecting $H$ and $\nu$ in Sec. III.1. We demonstrate in Sec. III.2 that the minimal TOs in both $mH^{(\nu)}_{\mathcal{D}}$ and $mH^{(\nu)}_{\rm GSD}$ are Abelian with $2q$ anyons when $q$ is odd, and $4q$ anyons when $q$ is even. The proof is relatively straightforward and illustrative of how we tackle the full $G$.
Having warmed up we dive into a discussion on constraints on the minimal TOs consistent with time reversal as well i.e., the full $G$. In Sec. III.3 we construct a fermionic Abelian TO which preserves $G$ and has a GSD of $4q^{2}$ ($q^{2}$) for $q$ even (odd). We then prove a number of results. First, in Sec. III.4, we show that any TO at filling $\nu=1/q$ which respects $G$ and has a GSD equal to or smaller than our construction must be Abelian. Moreover, we show that all such possibilities must have exactly the same GSD as our construction, so our construction is minimal. Then, in Sec. III.5, we show that our construction is the unique minimal TO that obey all the requirements. We defer a full classification of the symmetry enrichments of this minimal TO to future work.
@@
  • If our GSD is even smaller than this, then we can conclude that translation symmetry must be spontaneously broken. This will allow us to have a GSD as small as $q$ e.g., in a stripe order. Note that the GSD of $q$ in this case is not topological but arises from the breaking of translation symmetry.
@@ III.1 Generalities about TOs at fractional filling
In any gapped TO, bosonic or fermionic, with $\mathrm{U}(1)$ symmetry, there exists an anyon $v$ called a vison, [^10] The name ‘vison’ is used here as a generalization of its original use Senthil and Fisher (2000) to denote an Ising vortex in a $\mathbb{Z}_{2}$ gauge theory. In the special situation in which the low energy physics is described by a $\mathbb{Z}_{2}$ gauge theory with a charge-$1/2$ (under the global $\mathrm{U}(1)$) chargon, and a charge-neutral vison, the latter is indeed nucleated by such a flux insertion. which is nucleated by flux insertion, whose braiding with other anyons detects their (fractional) $\mathrm{U}(1)$ charge:
$\theta_{v,b}=2\pi Q_{b}\pmod{2\pi},$ (6)
where $Q_{b}$ is the charge carried by $b$. Physically, $v$ can be understood as resulting from the adiabatic insertion of a $2\pi$ $\mathrm{U}(1)$ flux. It is important to keep in mind that in a fermionic system with $\mathrm{U}(1)_{f}$ symmetry, all bosonic (fermionic) local excitations carry even (odd) charges. This means we can take $Q_{b}$ to be defined modulo two. [^11] The $\mathrm{U}(1)$ charge respects fusion, so $Q_{z}=Q_{x}+Q_{y}\pmod{2}$ for any $z$ in the fusion product of $x$ and $y$.
In addition, the classic Laughlin argument shows that $Q_{v}=\sigma_{H}$ (the Hall conductance in units where electronic charge and Planck’s constant are set to 1). It follows from Eq. (6) that $\theta_{v,v}=2\pi Q_{v}=2\pi\sigma_{H}$. The topological spin of $v$ must therefore be $\theta_{v}=\pi\sigma_{H}$ or $\pi\sigma_{H}+\pi$. If continuous rotation symmetry is assumed then it can be shown that $\theta_{v}=\pi\sigma_{H}$ Goldhaber et al. (1989); Greiter and Wilczek (2022). This can also be shown without continuous rotation symmetry Kapustin and Sopenko (2020), so we always have $\theta_{v}=\pi\sigma_{H}$.
@@
We note that the background anyon $a$ must be Abelian. This was shown more formally in Ref. Cheng et al. (2016), but we sketch two complementary arguments here. First suppose that $a$ is non-Abelian, with quantum dimension $d_{a}>1$. Then the Hilbert-space dimension associated with having non-Abelian anyons on every site will scale as $d_{a}^{N}$, where $N$ is the number of sites Kitaev (2006). This macroscopic degeneracy must be lifted by some interaction, meaning that this description is at best useful at an intermediate energy scale, and does not correspond to the true infrared (IR) limit. Flowing towards the true IR, any further macroscopic degeneracy will continue to be broken until we have fully broken it, at which point we must have a background anyon with $d_{a}=1$ i.e., $a$ must be Abelian.
Another complementary perspective can show the same conclusion. Operationally, the background anyon is defined by the following thought experiment: (1) excite some anyon pair $x$ and $\overline{x}$, (2) adiabatically transport $x$ around a unit cell, (3) and annihilate the pair. This will return the system to its ground state, so it can only lead to a global phase $\varphi_{x}$ for the wave function. We can think of this phase as the braiding of $x$ around the background anyon $a$. Importantly, this phase is additive under fusion: if $z\in x\times y$, then $\varphi_{z}=\varphi_{x}+\varphi_{y}$. Via a result in Ref. Barkeshli et al. (2019) [see Eq. (45)], this is then enough to establish that $\varphi_{x}=\theta_{a,x}$ for some Abelian anyon $a$, which we identify as the background anyon.
@@ III.2 Time-reversal broken minimal order
For bosons at $1/q$ filling, with $q$ even, $\mathrm{U}(1)_{q}$ (the TO of the bosonic $1/q$ Laughlin state) is a natural candidate for time-reversal breaking TOs in bosonic QCLs. We now demonstrate that for $q$ even, $\mathrm{U}(1)_{q}\boxtimes\{1,c\}$ cannot be a fermionic TO at filling $1/q$. First we note that for even $q$ $\mathrm{U}(1)_{q}$ by itself is a bosonic TO, and local operators in the theory are all bosons. It follows that the local excitation with minimal nonzero charge in $\mathrm{U}(1)_{q}$ is a charge-$2$ Cooper pair. Since every anyon in this theory has order $q$, the minimal fractional charge in this theory is $2/q$. But this means this theory does not have a background anyon of charge $Q_{a}=\nu=1/q$, contradicting the presence of $\mathrm{U}(1)_{f}$ symmetry and the filling constraint.
We can construct a possible fermionic TO by pairing two fermions into Cooper pairs. The Cooper pairs are bosons at a filling $\nu/2=1/2q$ per unit cell. We then make a TO, $\mathcal{C}$, which is translation invariant by considering $\mathrm{U}(1)_{2q}$, the bosonic Laughlin state of the Cooper pairs. Strictly speaking we have $\mathrm{U}(1)_{2q}\boxtimes\{1,c\}$, where $c$ is the fundamental fermion. This will be a fermionic TO in SETH,ν, and is depicted in Fig. LABEL:fig:cooper. This theory has a GSD of $2q$. Of course if $q$ is odd the theory $\mathrm{U}(1)_{q}$ is a fermionic TO in SETH,ν with a smaller GSD, $|\mathrm{U}(1)_{q}|/2=q$.
We now prove that these two constructions belong to $mH^{(\nu)}_{\rm GSD}$ for the respective choice of $q$. We do this by showing that the Abelian sector in any fermionic TO in SETH,ν is at least as large as our constructions. So then if our theory is to be minimal according to $mH^{(\nu)}_{\rm GSD}$ it must be Abelian with exactly the same anyon count as our constructions. This will prove that every TO in $mH^{(\nu)}_{\rm GSD}$ is Abelian and thus that $mH^{(\nu)}_{\rm GSD}=mH^{(\nu)}_{\mathcal{D}}$, so the two definitions of minimality coincide. Our proof strategy is illustrated in Fig. 4.
@@
Although we have determined the possible minimum TOs, we can still ask whether there are distinct symmetry enrichments by $\mathbb{Z}^{2}\times\mathrm{U}(1)_{f}$. To that end, we need to specify the vison anyon. It is easy to see that $v=a^{k}$, where $k$ is such that $2nk=1\pmod{q}$. Since $n$ is coprime with $q$ and $q$ is odd, such $k$ always exists and is uniquely fixed. Thus the possible minimal orders in the time reversal breaking case have Hall conductances of $\sigma_{H}=k/q\pmod{1}$ where $k$ can take all values relatively prime to $q$. An example of such a QCL phase was recently demonstrated in Chen et al. (2025). It is worth noting that we can (trivially) modify the TO by stacking copies of invertible topological states [^12] In two dimensions, the basic such state is the so-called $E_{8}$ state with chiral central charge $c=8$. of charge-neutral local bosonic excitations, which does not change the anyon content.
@@
Next we tackle $q$ even. In this case we again must have that $a^{q}\neq 1$. So then $a$ must have order a multiple of $2q$. Suppose now that $a^{q}=c$, then $a^{q}$ must braid trivially with $a$. This means that $\theta_{a,a^{q}}=2q\theta_{a}=0$. Since $q$ is even we multiply by the integer $q/2$ to see that $0=q^{2}\theta_{a}=\theta_{a^{q}}$ and thus $a^{q}$ is a boson, a contradiction with $a^{q}=c$. So $c\notin\{1,a,\ldots,a^{2q-1}\}$ and we must have that there are at least $4q$ anyons in our theory. For our theory to be minimal it must thus be Abelian with $\mathcal{A}=\langle a\rangle\boxtimes\{1,c\}$ where $a$ has order exactly $2q$. All such theories are again classified Moore and Seiberg (1989); Bonderson (2007), with $\theta_{a}=\pi n/2q$ for any $n$ such that $n$ is coprime with $q$. In particular this means $n$ must be odd. We can then choose $v=a^{k}$, where $k$ is such that $nk=2\pmod{2q}$. This means $k$ must be even, and more specifically that $k=2k^{\prime}$ with $k^{\prime}$ relatively prime to $2q$. Note $n=1$ corresponds to the $\mathrm{U}(1)_{2q}$ theory, where we can choose $v=a^{2}$. Thus when $q$ is even the minimal orders have Hall conductances $\sigma_{H}=2k^{\prime}/q\pmod{1}$, where $k^{\prime}$ can take all values relatively prime to $q$.
@@
This general proof strategy will prove useful in dealing with the full $G$. We first construct a theory in SETG,ν, then show that all minimal theories must be at least as big as our construction. We again refer to reader to Fig. 4 for an illustration of this strategy. Having obtained the anyon count of each minimal theory we then show that in the presence of time-reversal symmetry the minimal TO is in fact unique.
@@
Figure 4: Our proof strategy for demonstrating the number of anyons a minimal theory must have. Using Cooper pairing we construct a theory in SETG,ν. This theory’s anyon count must upper bound the anyon count of any minimal SET. We further consider all possible powers of the Abelian anyons we know our theory must possess. This will lower-bound the anyon count. Any putative minimal SET will be caught in the vice between these two bounds.